If Rcosx = 1 and Rsinx = 3, where R>0 and 0 I know it has something to do with further Trig equations for year 11 3 unit maths.
Answer by Alan3354(69443)

If Rcosx = 1 and Rsinx = 3, where R>0 and 0 < x < 90, find the exact value of R and x correct to the nearest degree
------------------
sin = 3*cos
sin^2 + cos^2 = 1
9cos^2 + cos^2 = 1
cos^2 = 1/10
cos(x) = sqrt(10)/10
x =~ 71.565 degs --> 72 degs to the nearest deg
------
R*cos(x) = 1
R = 1/cos(x) = sqrt(10)